OpenStudy (anonymous):

is the slope -1 for (0,3) and (-6,-4) I been trying this slope thing and I hope I finally got it

OpenStudy (anonymous):

\(Slope = \frac{\Delta y}{\Delta x}\) \( = \frac{y_1-y_0}{x_1-x_0} = \frac{3-(-4)}{0-(-6)} = \frac{3+4}{0+6} = \frac{7}{6}\)

OpenStudy (anonymous):

thanks